# Improving Visual Place Recognition with Sequence-Matching Receptiveness Prediction

### v1

@@ V-B Ablation Study: Varying Sequence Length
We analyze the influence of sequence length in sequence matching on our SMR predictor. Figure 3a shows the results for all seven VPR models evaluated on the Nordland dataset. As in Section V-A, we use our predictor to remove matches predicted as incorrect after sequence matching in the VPR system with sequence matching. The bottom of each bar represents the PR AUC of the VPR system with sequence matching, while the top indicates the PR AUC when our predictor is applied to remove incorrect matches. If the difference is positive, the bar is colored green (as is the case for all results in Figure 3a); otherwise, it would have been red.
@@ V-C Ablation Study: Varying Predictor’s Trust (First Filter)
We evaluate how varying the trust level in our SMR predictor affects the precision-recall plot of the VPR system with sequence matching (4-frame sequence length).
Figure 4 shows the precision-recall plot of our SMR predictor at different trust levels for match removal. When we only trust the predictor at full confidence (), very few to no matches are removed. In this case, the effect of our predictor is minimal, and the precision-recall plot closely mirrors that of the VPR system with sequence matching.
At the other end of the spectrum, fully trusting our predictor whenever it predicts a match as incorrect after sequence matching – corresponding to a in this case (EigenPlaces on Nordland) – results in the highest precision improvement at lower recall values compared to other trust levels. Since the predictor removes matches predicted as incorrect after sequence matching, the resulting precision-recall curve does not extend to a recall of . In Section V-F, we examine how our predictor performs when it is first used to remove incorrect matches after sequence matching and then used as a second filter to replace the removed match with another highly ranked match.
@@ V-D Ablation Study: Impact of Each Attribute
In Table II, we performed an ablation study on the Nordland dataset to assess the impact of each attribute individually on the F1-score of our MLP-based SMR predictor in isolation, i.e., without applying it to the VPR system with sequence matching. Results are averaged across all seven VPR models. Using only A1 or A3 resulted in an F1-score of , while using only A2 or A4 resulted in an F1-score of . Using all four attributes improved performance to . This demonstrates that incorporating all four attributes improves classifier performance, validating our approach.
@@ V-E Qualitative Results
We present qualitative results in Figure 5, showing examples where the baseline VPR system with sequence matching produces incorrect predictions. We show cases where our predictor effectively removes incorrect matches and restores them with another highly-ranked match. We also highlight instances where the predictor fails to detect or restore incorrect matches.
@@ V-F Match Restoration
Can discarded matches be replaced by alternative correct matches? This section presents the results for applying our predictor a second time, this time to rank the alternative highly ranked matches for a discarded match in order to replace it with the most credible alternative among the next top three ranked matches.
@@ VI Discussion
We presented a novel approach for predicting the receptiveness of single-frame VPR methods to distance-based sequence matching, which is independent of the underlying VPR model. By removing matches output by the sequence matching system that are predicted to be incorrect, we can substantially improve the average VPR performance of the underlying VPR + Sequence Matching system, a result demonstrated across seven VPR models and three benchmark datasets. Importantly, the resulting performance improvement is superior to simply choosing a different operating point on the baseline system’s precision-recall curve – that is, performance after applying the predictor moves above the baseline precision-recall curve in most cases. While the predictor mechanism used in this paper, an MLP, is relatively straightforward, it is particularly interesting to note that even with imperfect prediction performance, such a substantial average improvement in VPR performance can be achieved.
@@
We believe this work is a promising expansion on the recent single-frame introspective work (covered in Section II-C) to introspection on filtering and sequence-based methods, which are widely used. Intuitively, high-performance introspection may be more achievable on sequential information over time, as there is more information available, and in robotics applications, there are interesting correlations across sequences in this information. Some of the most fruitful areas of future exploration would be investigating the utility of modern machine learning techniques with particular mechanistic analogues to this process – like transformers.
@@
While in this work, significant average VPR performance improvement was achievable by applying the predictor to remove some matches that are predicted to be incorrect, there is great appeal in potentially restoring lower ranked candidate matches. We investigated this possibility in this paper, showing that match restoration appears to be a fundamentally more difficult task, at least in the context of providing further performance improvements over the main predictor-filter approach demonstrated in this work. More sophisticated predictor learning schemes may be able to obtain the predictor performance required to achieve further VPR performance improvements: in parallel simulations (not presented in this paper) we have discovered that the predictor performance required to beneficially restore matches is much higher than to filter out bad matches.
@@
Computational efficiency was not the primary aim of the research presented here, although Section V-G presents a brief description of the inference computation that shows this is a very lightweight process. The current system is a post-sequence-matching process, in that it runs after sequence matching has been performed. Future work could investigate the possibility of using our predictor in an incremental manner (e.g., using shorter sequences and only on a subset of the data) to guide sequence matching before it is fully computed, enabling both potential computational benefits and matching latency reductions, with the predictor being applied live, e.g., to determine whether increasing the sequence length at that moment in time is worthwhile.

### v2

@@ V-B Ablation Study: Varying Sequence Length
We analyze the influence of sequence length in sequence matching on our SMR predictor. Figure 3a shows the results for all seven VPR models evaluated on the Nordland dataset. As in Section V-A, we use our predictor to remove matches predicted as incorrect after sequence matching in the VPR system with sequence matching. The bottom of each bar represents the PR AUC of the VPR system with sequence matching, while the top indicates the PR AUC when our predictor is applied to remove incorrect matches. If the difference is positive, the bar is green (as in all results in Figure 3a); otherwise, it would be red.
@@ V-C Ablation Study: Varying Predictor’s Trust (First Filter)
We evaluate how varying the trust level in our SMR predictor affects the precision-recall plot of the VPR system with sequence matching (4-frame sequence length). Figure 4 shows the PR plot of our SMR predictor at different trust levels for match removal on EigenPlaces and NetVLAD. When we only trust the predictor at full confidence (), very few to no matches are removed. In this case, the effect of our predictor is minimal, and the PR plot closely mirrors that of the VPR system with sequence matching.
For NetVLAD, lowering the trust threshold allows the predictor to remove more incorrect matches after sequence matching, resulting in a clear gain in precision at lower recall values. In contrast, for EigenPlaces, the benefit is marginal across trust levels, with the precision-recall curves closely placed. This suggests that in a high-performing VPR model (like EigenPlaces) with sequence matching, it is more challenging to remove the few remaining incorrect matches, as they are already sparse, while effectively retaining the correct matches. As the predictor removes matches predicted as incorrect after sequence matching, the resulting PR curve does not extend to a recall of . Section V-F examines how our predictor performs when it is first used to remove incorrect matches after sequence matching and then used as a second filter to replace the removed match with another highly ranked match.
@@ V-D Ablation Study: Impact of Each Attribute
In Table II, we performed an ablation study on the Nordland dataset to assess the impact of each attribute individually on the F1-score of our MLP-based SMR predictor in isolation, i.e. without applying it to the VPR system with sequence matching. Results are averaged across all seven VPR models on the Nordland dataset. Using only A1 or A2 resulted in an F1-score of , while using A3 and A4 resulted in F1-scores of and , respectively.
Using all four attributes improved performance to . This demonstrates that incorporating all four attributes improves classifier performance, validating our approach. These attributes were selected through an iterative process of visualizing distance scores, analyzing top-K matches, deriving descriptive statistics, and examining class-wise sample distributions.
@@ V-E Qualitative Results
We present qualitative results in Figure 5, showing examples where the baseline VPR system with sequence matching produces incorrect predictions. We show cases where our predictor effectively removes incorrect matches and restores them with another highly-ranked match, and instances where the predictor fails to detect or restore incorrect matches.
@@ V-F Match Restoration
Can discarded matches be replaced by alternative correct matches? This section presents the results for applying our predictor a second time, this time to rank the alternative highly ranked matches for a discarded match to replace it with the most credible alternative among the next top three ranked matches.
@@ VI Discussion
We presented a novel approach for predicting the receptiveness of single-frame VPR methods to distance-based sequence matching, which is independent of the underlying VPR model. By removing matches output by the sequence matching system that are predicted as incorrect, we can substantially improve the average VPR performance of the underlying VPR + Sequence Matching system, a result demonstrated across seven VPR models and three benchmark datasets. Importantly, the resulting performance improvement is superior to simply choosing a different operating point on the baseline system’s precision-recall curve – that is, performance after applying the predictor moves above the baseline precision-recall curve in most cases. While the predictor mechanism used in this paper, an MLP, is relatively straightforward, it is particularly interesting to note that even with imperfect prediction performance, such a substantial average improvement in VPR performance can be achieved.
@@
We believe this work is a promising expansion on the recent single-frame introspective work (covered in Section II-C) to introspection on widely used filtering and sequence-based methods. Intuitively, high-performance introspection may be more achievable on sequential information over time, as there is more information available, and in robotics applications, there are interesting correlations across sequences in this temporal data. Some areas for future exploration would be investigating the use of modern machine learning techniques with particular mechanistic analogues to this process, such as Transformers.
@@
While in this work significant average VPR performance improvement was achievable by applying the predictor to remove some matches that are predicted to be incorrect, there is great appeal in potentially restoring lower ranked candidate matches. We investigated this possibility in this paper, showing that match restoration appears to be a fundamentally more difficult task, at least in the context of providing further performance improvements over the main predictor-filter approach demonstrated in this work. More sophisticated predictor learning schemes may be able to obtain the predictor performance required to achieve further VPR performance improvements: in parallel simulations (not presented in this paper) we have discovered that the predictor performance required to beneficially restore matches is much higher than to filter out bad matches.
@@
Computational efficiency was not the primary aim of the research presented here, although Section V-G presents a brief description of the inference computation that shows this is a very lightweight process. The current system is a post-sequence-matching process, in that it runs after sequence matching has been performed. Future work could investigate the possibility of using our predictor in an incremental manner (e.g. using shorter sequences and only on a subset of the data) to guide sequence matching before it is fully computed, enabling both potential computational benefits and matching latency reductions, with the predictor being applied live, e.g. to determine whether increasing the sequence length at that moment in time is worthwhile.